the length of a rectangular swimming pool is 12 ft less than its length. If the perimeter of the pool is 96 ft, what are the dimensions of the pool?
Answer by checkley77(12844)

P=2L+2W
96=2L+2(L-12)
96=2L+2L-24
4L=96+24
4L=120
L=120/4
L=30 FT.
W=30-12=18 FT.
PROOF
96=2*30+2*18
96=60+36
96=96